OpenStudy (anonymous):

I will give out a medal and I'll fan you! A company distributes free crayon boxes to all the students of x schools. Each school has (x + 5) classes. The number of students in each class is 2 less than the number of classes in each school. Each student is given 2 boxes of crayons. Part A: Write an expression to show the total number of crayon boxes distributed by the company in x schools. (4 points)

OpenStudy (anonymous):

Number of schools = x Number of classes in each school = (x + 5) Number of students in each class = Number of classes in each school - 2 = x + 5 - 2 = (x + 3) Number of crayons for each student Total number of crayons distributed = Number of crayons for each student * Number of students in each class * Number of classes in each school * Number of schools

OpenStudy (anonymous):

I'm sorry to say this, but @ranga posted that, and it wasn't very helpful. :/

OpenStudy (anonymous):

May I please have some help? I'm not very good at this. :')

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@reaper123 any luck?

OpenStudy (anonymous):

i knew thats who posted it lol , i was just trying to help. sorry xc

OpenStudy (anonymous):

It's ok.

OpenStudy (anonymous):

Do you understand what it means though?